Ethanol (C2H5OH) has a measured density of 0.789 g/mL. You need a solution of ethanol with a concentration of 0.8875 M. 

The final volume of the solution should be 0.545 L. What volume of ethanol needs to be added to the flask before it is filled with water to its final volume?


1
Expert's answer
2021-02-28T06:20:24-0500

C=n/V

n=C*V=0.8875*0.545=0.4837(mol)

m=n*M

M=2*12+6*1+16=46(g/mol)

m=0.4837*46=22.2502(g)

V=m/d=22.2502/0.789=28.2(ml)